5.Find the percentage by mass of each of the elements in C6H5Cl. Round the answers to the nearest 0.01%
Answer #1

% by mass of an element = (mass of the element/ total mass of compound) x 100

% by mass of C = (12.01x6 / 112.5) x 100 = 64.03%

% by mass of H = ( 1.008x5 / 112.5)x100 = 4.48%

% by mass of Cl = ( 35.5/112.5)x100 = 31.56%